OpenStudy (surryyy):

The area of a square is 50 square meters. How long is each side? Round your answer to the nearest meter. Answer 2,500 100 50 7

OpenStudy (jhannybean):



OpenStudy (jhannybean):

\[A_s = l \cdot w = x\cdot x = x^2\]

OpenStudy (jhannybean):

So if they give you area ,\(A=50 \ m^2\), then \[50 \ m^2 = x^2\]\[x = \sqrt{50} = ?\]